### v2

# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

Abstract
Recent multi-LLM agent systems have shown promising capabilities for prompt optimization and automated problem solving, but prior approaches rely on frozen fine-tuned agents or a single fine-tuning pipeline. We propose an ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution) model in which multiple supporter agents collaboratively train and refine a self-evolving agent toward a domain-specific research agent. Supporter agents specialize in complementary roles, including exploration support, fine-tuning strategy, and adaptive reference-policy inspection for the proposed Evolving Direct Preference Optimization (EvoDPO). Within this framework, we introduce a diverse set of challenging problem instances and demonstrate that our method achieves substantial performance improvements over existing approaches across long horizons.
Machine Learning, ICML
1 Introduction
Fine-tuned large language model (LLM) agents are broadly used as components in agentic systems for complex problem solving, scientific computing, and code-generation workflows. Recent frameworks show that collections of specialized agents, combined with experience banks, routing, or collaborative coordination, improve task performance and reliability by optimizing search, planning, decision-making, strategies, or data generation (Liu et al., 2025; Jiang and Karniadakis, 2025; Wu et al., 2025; He et al., 2025; Chen et al., 2025). In parallel, multi-agent infrastructure and benchmarking efforts have made it easier to orchestrate such systems and study their behaviors at scale (Chen et al., 2023; Wu et al., 2023; Chen et al., 2024; Huang et al., 2025; Becker et al., 2025).
Despite these advances, many existing approaches predominantly treat LLM agents as frozen optimizers and focus on improving optimization efficiency rather than explicitly studying how to develop an LLM agent through iterative model updates under realistic non-stationary and long-horizon training. To bridge this gap, our work introduces a framework for robust self-evolution via controlled reference management.
In this work, we propose ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution), a framework targeting specialized research agent formation via systematic model development. Rather than relying on a single static agent, we fine-tune a set of supporter LLM agents that collectively cultivate and refine a domain-specific research agent. The supporter agents are assigned complementary roles that support the learning dynamics: (i) exploration support to suggest diverse exploration strategies (ii) fine-tuning supervision to stabilize preference-based updates and prevent excessive deviation, and (iii) reference-policy inspection to decide when and how to renew the reference policy for well-aligned adaptation. By explicitly separating these tasks, the training loop prioritizes progressive agent improvement rather than merely search or planning enhancements, and it remains compatible with scalable multi-agent execution (Xie et al., 2024).
Our learning mechanism leverages preference optimization as a practical alternative to Reinforcement Learning from Human Feedback (RLHF) pipelines. Direct Preference Optimization (DPO) enables preference learning without an explicit reward model (Rafailov et al., 2024), and recent work extends preference learning beyond simple binary comparison and introduces progressive schemes to handle distribution shift (Liao et al., 2025; Yang et al., 2025). Our work aligns with evolutionary or multi-agent fine-tuning approaches that iteratively improve a developing agent using critique or debate signals (Surina et al., 2025; Zhou et al., 2025). However, a key limitation in many iterative preference-based pipelines is fixed reference policies, which leads to misaligned references, overly conservative updates, or stagnation.
To address this, we introduce Evolving DPO (EvoDPO), a preference-optimization loop with telemetry-driven fine-tuning control and adaptive reference management. At each fine-tuning phase, a strategist agent tunes DPO hyperparameters based on training diagnostics. In parallel, EvoDPO updates its policy using the DPO algorithm, while modeling the reference policy as a phase-indexed variable. Specifically, the self-evolving agent proposes a Kullback-Leibler (KL) regularized reference policy to be used in the next phase. Then, a policy-inspector agent applies a KL trust-region gate to decide whether to promote the proposed policy as the next reference or to keep the existing one. This integration of adaptive preference optimization and task-distributed agents prevents stagnation from stale references while avoiding instability from overly aggressive reference adaptation, enabling robust long-horizon self-evolution. The resulting mechanism is complementary to recent multi-agent preference-alignment approaches (Lyu et al., 2025).
To rigorously evaluate our core algorithm, EvoDPO, we conduct a comparative analysis against a fixed-reference DPO baseline within a theoretical framework. Specifically, we formulate a non-stationary contextual bandit problem and analyze the dynamic regret to formally characterize and demonstrate the performance advantages of our algorithm.
In the numerical results section, we compare our proposed methods against EvoTune, a single-agent evolution baseline that performs isolated evolutionary exploration and fixed-reference fine-tuning (Surina et al., 2025). We evaluate our framework on two complementary settings. First, we consider contextual multi-armed bandits under non-stationary environments (Wu et al., 2018), providing a controlled setting for adaptive policy evolution under the drifting model parameter and connecting to recent LLM-assisted bandit initialization ideas (Alamdari et al., 2024; Toscano et al., 2025). Second, to demonstrate applicability to scientific machine learning (SciML), we address the 1D Burgers’ equation using physics-informed neural networks (PINNs). Instead of relying on fixed, hand-tuned loss weights, ATLAS adaptively reweights PINN loss terms, effectively designing the training objective during optimization. This is particularly relevant in challenging regimes, such as those exhibiting near finite-time blow-up behavior (Kumar and Mukherjee, 2024). Overall, these evaluations indicate that our framework can sustain iterative improvement in both decision-making and SciML settings.
In this work, our contributions are as follows:
(i) Task-distributed multi-LLM agents structure for self-evolving research agent formation
(ii) Adaptive reference management mechanism for the EvoDPO algorithm
(iii) Numerical results showing optimal problem solving under long-horizon self-improvement
2 Proposed Methods
2.1 ATLAS : Adaptive Task-distributed Learning for Agentic Self-evolution
ATLAS is the end-to-end self-evolution framework comprising (i) a task-distributed multi-LLM supporter layer, (ii) a preference data construction procedure, and (iii) an adaptive fine-tuning loop for a developing research agent. Figure 1 illustrates the overall architecture of our framework and the specific workflow of each separated agent.
Self-evolving research agent.
Before each fine-tuning stage, the research agent performs a six-island exploration procedure in parallel to generate a diverse set of candidates (Surina et al., 2025). Each island maintains an independent candidate population and an update rule that mitigates mode collapse and encourages specialization across islands. For each explored candidate, the agent stores the task score and auxiliary metadata in an island-specific buffer. To preserve diversity, the agent also maintains a clustering over buffered candidates and tracks cluster coverage during selection. In particular, the agent monitors the number of clusters represented both in the buffer and within the count-based top-$s$ candidates, and the agent prioritizes candidates that preserve diverse coverage rather than concentrating on a single mode. At each exploration round, each island ranks candidates by score and uses the top-$s$ to form preference pairs. This procedure biases exploration toward high-scoring regions while preserving diversity via the clustering signal.
Exploration is implemented through a producer-consumer architecture. The producer assigns prompts to each island and the consumers evaluate candidates to produce task scores and metadata. The system also selects the inference model used by each island, enabling consistent generation behavior within an island while allowing diversity across islands.
For preference learning, the agent partitions evaluated candidates into passed and failed sets according to the task evaluation. It then selects count-based top-$s$ passed candidates to form preferred and less-preferred pairs $(y^{+},y^{-})$, and attaches auxiliary metadata $m$ for filtering and analysis. This process yields a specific preference dataset $D_{k}$
$D_{k}=\{(x_{i},y_{i}^{+},y_{i}^{-},m_{i})\}_{i=1}^{N_{k}},$
where $x$ denotes the prompt and $N_{k}$ is the number of preference pairs collected in phase $k$.
Figure 1: ATLAS workflow. ATLAS (Adaptive Task-distributed Learning for Agentic Self-evolution) alternates between (i) exploration with a supporter agent to generate diverse candidates and a preference dataset, and (ii) EvoDPO updates consisting of DPO fine-tuning (strategist-guided) and reference promotion via an inspector gate based on score improvement and a KL budget.
Let $\pi_{\theta,k}$ denote the current research agent policy and $\pi_{ref,k}$ the reference policy used at fine-tuning phase $k$. After constructing $D_{k}$, the research agent updates the policy by minimizing the DPO objective, and proposes a KL-regularized reference for the next phase $k+1$.
Task-Distributed Multi-LLM Supporters.
In our multi-LLM agentic system, the framework consists of (i) a developing Research Agent that is iteratively updated, and (ii) a set of frozen, fine-tuned supporter LLM agents specialized for distinct roles. We initialize the developing research agent from a general pre-trained model, whereas the supporter agents start from instruction-tuned or domain-adapted checkpoints. This asymmetric initialization enables the supporter agents to provide strong guidance and critique, particularly in early training.
Table 1: LLM agents and their assigned roles.
Model Role / Task allocation
gpt-oss-120B Exploration Supporter
DeepSeek-R1-32B Fine-Tuning Strategist
Llama4-latest Policy Inspector
Llama-3.2-1B Research Agent
As shown in Table 1, we employ three supporter agents: (i) an Exploration Supporter that proposes diverse candidate completions and provides exploration strategy suggestions, (ii) a Fine-Tuning Strategist that recommends fine-tuning hyperparameters (e.g., inverse-temperature $\beta$, batch size, and learning rate) conditioned on training diagnostics, and (iii) a Policy Inspector that performs adaptive reference management by deciding whether to update the reference policy under a conservative safety constraint.
Algorithm 1 ATLAS
1:Initialize: island buffer $\{\mathcal{P}_{i}\}_{i=0}^{5}$, research agent policy $\pi_{\theta,0}$, reference $\pi_{ref,0}\leftarrow\pi_{\theta,0}$
2:Set: fine-tuning round $k\leftarrow 0$
3:for $t=1,2,\dots$ do
4:  for island $i\in\{0,\dots,5\}$ in parallel do
5:   Sample prompt $x_{i,t}\sim\mathcal{X}$
6:   Candidate set $\mathcal{C}_{i,t}=\{y_{i,t}^{(j)}\}_{j=1}^{J}\sim\pi_{\theta,k}(\cdot\mid x_{i,t})$
7:   for $j=1,\dots,J$ do
8:    Evaluate $y_{i,t}^{(j)}$ to obtain score $s_{i,t}^{(j)}$ and metadata $m_{i,t}^{(j)}$
9:    $\mathcal{P}_{i}\leftarrow\mathcal{P}_{i}\cup\{(x_{i,t},y_{i,t}^{(j)},s_{i,t}^{(j)},m_{i,t}^{(j)})\}$
10:   end for
11:  end for
12:  if fine-tuning flag = 1 and $k<K$ then
13:   Construct preference dataset $D_{k}$ from $\cup_{i}\mathcal{P}_{i}$
14:   Exploration Supporter: proposes strategies
15:   $D_{k}\leftarrow D_{k-1}$
16:   for selected prompts $x$ do
17:    Form a preference pair $(y^{+},y^{-})$ for $x$
18:    $D_{k}\leftarrow D_{k}\cup\{(x,y^{+},y^{-},m)\}$
19:   end for
20:   Execute Algorithm 2 (EvoDPO)
21:    Fine-Tuning Strategist: Set hyperparameters
22:    Policy Inspector: Reference management
23:   Set $k\leftarrow k+1$
24:  end if
25:end for
At the fine-tuning phase $k$, the developing research agent executes an adaptation loop. Given prompts $x_{t}\sim\mathcal{X}$ and the current research agent policy $\pi_{\theta,k}$, the Exploration Supporter is invoked during the fine-tuning phase to construct preference supervision.
For each prompt $x_{t}$, it proposes a diverse set of candidate completions, targeted perturbations, and exploration parameter-tuning strategies, conditioned on recent failures and task constraints.
In parallel, a Fine-Tuning Strategist consumes training telemetry from the current stage (e.g., preference-set size, score distributions, and recent accept/reject events) and returns a control signal for the next stage. These signals update hyperparameters such as $\beta$, learning-rate schedules, pair-selection thresholds, and regularization strengths.
Finally, the Policy Inspector performs gated reference management. It evaluates the proposed KL-regularized policy $\pi^{kl}_{k}$ relative to current reference $\pi_{ref,k}$ and decides whether to set the reference to $\pi_{ref,k+1}\leftarrow\pi^{kl}_{k}$ or keep $\pi_{ref,k}$, subject to a conservative safety constraint that prevents aggressive drift. We detail these mechanisms in section 2.2.
This separation of responsibilities stabilizes exploration, adapts the optimization dynamics, and prevents uncontrolled reference drift over a long-horizon. Within ATLAS, our core algorithm, EvoDPO, performs a progressive and robust adaptation.
2.2 Evolving DPO algorithm
In this section, we introduce our core algorithm, Evolving DPO (EvoDPO), that performs preference optimization with KL-regularized reference improvement.
Direct Preference Optimization.
Direct preference optimization (DPO) (Rafailov et al., 2024) is an efficient alternative to reward-model-based RLHF. It optimizes a policy from pairwise preferences and avoids explicit reward learning and value-function estimation. Given a preference dataset $D$ of $(x,y^{+},y^{-})$, DPO minimizes
$L_{\mathrm{DPO}}$ $=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}\Biggl[\log\sigma\Biggl(\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\mathrm{ref}}(y^{+}\mid x)}$ (1)
$\qquad\qquad-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\mathrm{ref}}(y^{-}\mid x)}\Biggr)\Biggr]$
where $\pi_{\mathrm{ref}}$ is a fixed reference policy and $\beta>0$ controls the strength of the preference margin.
A practical assumption behind (1) is that the preference data distribution is reasonably aligned with the reference policy: many public preference datasets are collected from, or near, a supervised fine-tuning (SFT) policy, and optimization is most stable when the likelihood ratios in (1) are not evaluated far off-distribution. In long-horizon phase-indexed loops, however, a fixed $\pi_{\mathrm{ref}}$ can become outdated as the research agent improves its policy and dataset. In this case, the mismatch between $D_{k}$ and $\pi_{\mathrm{ref}}$ can yield slow adaptation and misalignment.
Evolving DPO.
To address reference-data mismatch in long-horizon self-improvement, we introduce Evolving DPO (EvoDPO), which replaces the fixed reference with a phase-indexed reference $\pi_{\mathrm{ref},k}$ that updates across phases. At phase $k$, the research agent is updated by optimizing a DPO objective anchored at the current reference:
$L_{\mathrm{EvoDPO}}$ $=-\mathbb{E}_{(x,y^{+},y^{-})\sim D_{k}}\Biggl[\log\sigma\Biggl(\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\mathrm{ref},k}(y^{+}\mid x)}$ (2)
$\qquad\qquad-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\mathrm{ref},k}(y^{-}\mid x)}\Biggr)\Biggr]$
yielding an updated research agent policy $\pi_{\theta,k+1}$. The reference is then updated conservatively by the reference management.
Empirical KL estimate in the gate.
The exact prompt-conditional divergence $D_{KL}$
$D_{KL}\big(\pi(\cdot\mid x)\,\|\,\pi_{\mathrm{ref},k}(\cdot\mid x)\big)$
between sequence-valued policies is intractable to compute, since it requires marginalization over the space of output sequences. We therefore use a tractable, trajectory-conditioned KL estimator that compares next-token distributions along a representative candidate trajectory while conditioning on the same prompt. Specifically, for each prompt $x\in G_{k}$, we associate a candidate response $y=(y_{1},\ldots,y_{L_{y}})$ from the phase buffer. In our implementation, $y$ is the preferred candidate $y^{+}$ associated with $x$. For each token position $\ell=1,\ldots,L_{y}$, define the history-conditional next-token distributions
$p_{x,y,\ell}(\cdot):=\pi(\cdot\mid x,y_{<\ell}),\quad q_{x,y,\ell}(\cdot):=\pi_{\mathrm{ref},k}(\cdot\mid x,y_{<\ell}),$
where $y_{<\ell}=(y_{1},\ldots,y_{\ell-1})$ denotes the prefix up to (but excluding) position $\ell$. We define the per-pair token-average KL as
$\bar{D}(x,y):=\frac{1}{L_{y}}\sum_{\ell=1}^{L_{y}}D_{KL}\bigl(p_{x,y,\ell}\,\|\,q_{x,y,\ell}\bigr),$
and the empirical gate statistic as
$\widehat{\mathrm{KL}}_{k}\bigl(\pi\,\|\,\pi_{\mathrm{ref},k}\bigr):=\frac{1}{|G_{k}|}\sum_{(x,y)\in G_{k}}\bar{D}(x,y).$ (3)
In implementation, $D_{KL}(p_{x,y,\ell}\|q_{x,y,\ell})$ is computed directly from model logits via softmax with standard causal alignment, and masking excludes the prompt tokens and averages only over completion tokens.
Inspector score functional.
At phase $k$, the Policy Inspector decides whether to promote the reference using a randomly sampled gate subset $G_{k}$ drawn from the current phase buffer. Each element of $G_{k}$ is a prompt $x$ together with its associated evaluated candidates and task scores collected during exploration. This design aligns the gate with the on-distribution prompts encountered by the developing agent during self-evolution, while reducing selection adaptivity relative to score-based selection.
Given a policy $\pi$, let $\mathrm{Score}(x,\pi)$ denote the task-specific evaluation score used by the inspector on prompt $x$. We define the inspector score as the subset average
$\widehat{\mathcal{S}}_{k}(\pi):=\frac{1}{|G_{k}|}\sum_{x\in G_{k}}\mathrm{Score}(x,\pi).$
The inspector uses the score difference relative to the current reference:
$\Delta\widehat{\mathcal{S}}_{k}(\pi):=\widehat{\mathcal{S}}_{k}(\pi)-\widehat{\mathcal{S}}_{k}(\pi_{\mathrm{ref},k}),\qquad\Delta\widehat{\mathcal{S}}_{k}:=\Delta\widehat{\mathcal{S}}_{k}(\pi^{kl}_{k}).$
Reference-improvement operator.
In practice, we approximate the ideal reference-improvement operator by optimizing over a finite candidate set $\mathcal{C}_{k}\subset\Pi$. At phase $k$, we construct $\mathcal{C}_{k}$ from (i) a small number of intermediate checkpoints of the developing agent collected during the phase, and (ii) candidate policies produced under the exploration and guidance from the Fine-Tuning Strategist. Thus, $\mathcal{C}_{k}=\{\pi^{(1)}_{k},\ldots,\pi^{(M)}_{k}\}$, and $\mathcal{C}_{k}$ always includes the phase-final policy $\pi_{\theta,k+1}$. With this candidate set, we define the practical reference-improvement operator as
$\pi^{kl}_{k}:=\arg\max_{\pi\in\mathcal{C}_{k}}\Bigl\{\widehat{\mathcal{S}}_{k}(\pi)-\beta_{ref}\,\widehat{\mathrm{KL}}_{k}\bigl(\pi\,\|\,\pi_{\mathrm{ref},k}\bigr)\Bigr\},$ (4)
where $\widehat{\mathcal{S}}_{k}$ and $\widehat{\mathrm{KL}}_{k}$ are computed using the phase-$k$ gate subset. Here, $\beta_{ref}>0$ is a reference-gating trade-off coefficient that controls how strongly the inspector penalizes reference drift during selection.
Algorithm 2 EvoDPO
1:Initialize: Research agent policy $\pi_{\theta,1}$, Reference $\pi_{\text{ref},1}\leftarrow\pi_{\theta,1}$
2:Hyperparameters: DPO temperature $\beta$, KL budget $\delta_{H}$, score tolerance $\epsilon_{s}$
3:for $k=1,\dots,K$ do
4:  Collect $D_{k}=\{(x_{i},y^{+}_{i},y^{-}_{i},m_{i})\}^{N_{k}}_{i=1}$
5:  Update research agent policy $\pi_{\theta,k}$$\pi_{\theta,k+1}\leftarrow\arg\min_{\pi\in\Pi}\mathcal{L}_{\text{DPO}}(\pi;\pi_{\text{ref},k},\mathcal{D}_{k})$
6:  Compute $\pi^{kl}_{k}$ via Eq. (4)
7:  Compute $\widehat{\mathrm{KL}}_{k}(\pi^{kl}_{k}\|\pi_{\mathrm{ref},k})$
8:  if $\Delta\widehat{\mathcal{S}}_{k}\geq\epsilon_{s}$ and $\widehat{\mathrm{KL}}_{k}(\pi^{kl}_{k}\|\pi_{\mathrm{ref},k})\leq\delta_{H}$ then
9:   Accept $\pi_{\text{ref},k+1}\leftarrow\pi^{kl}_{k}$
10:  else
11:   Reject $\pi_{\text{ref},k+1}\leftarrow\pi_{\text{ref},k}$
12:  end if
13:end for
Reference update with conservative acceptance.
At phase $k$, the research agent proposes the candidate reference $\pi^{kl}_{k}$ via Eq.  (4). Then the Policy Inspector applies a gating rule using (i) a score improvement $\Delta\widehat{\mathcal{S}}_{k}$ and (ii) a KL trust-region measured by the trajectory-conditioned token-level KL estimator $\widehat{\mathrm{KL}}_{k}$ on the phase-$k$ gate subset $G_{k}$. The system accepts a reference promotion if
$\Delta\widehat{\mathcal{S}}_{k}\geq\epsilon_{s}\quad\text{and}\quad\widehat{\mathrm{KL}}_{k}\bigl(\pi^{kl}_{k}\,\|\,\pi_{\mathrm{ref},k}\bigr)\leq\delta_{H},$ (5)
where $\epsilon_{s}\geq 0$ is a small tolerance that prevents noise-driven promotions and $\delta_{H}$ prevents aggressive shifts. If Eq.  (5) holds, the inspector updates the reference for the next phase, $\pi_{\mathrm{ref},k+1}\leftarrow\pi^{kl}_{k}$; otherwise, it maintains $\pi_{\mathrm{ref},k+1}\leftarrow\pi_{\mathrm{ref},k}$. This accept/reject rule is a conservative realization of the ideal KL-regularized progression in Eq.  (12), made operational by restricting to $\mathcal{C}_{k}$ in Eq.  (4) and using the $\widehat{KL}_{k}$.
2.3 Theorem
Theory–practice connection.
We provide a regret analysis for an idealized KL-regularized reference-update Eq. (12) that abstracts EvoDPO’s reference progression mechanism. The accept/reject inspector gate is a practical implementation-level safeguard and is empirically evaluated in section 3.
Problem Formulation
We consider a non-stationary preference-based contextual bandit that abstracts the reference-progression mechanism used in EvoDPO. At each round t, the environment reveals a context $x_{t}\in\mathcal{X}$, the agent samples two actions $y^{(1)}_{t},y^{(2)}_{t}\sim\pi_{t}(\cdot|x_{t})$, where the action space $\mathcal{Y}$ is a finite set. Thus $\pi_{t}(\cdot|x_{t})$ is a categorical distribution over $\mathcal{Y}$. The preference outcome follows a Bradley-Terry model with utility
$u_{t}(x,y):=\langle\theta_{t},\phi(x,y)\rangle$ (6)
$\Pr\!\left(y\succ y^{\prime}\mid x,t\right)=\sigma\!\left(u_{t}(x,y)-u_{t}(x,y^{\prime})\right)$ (7)
where $\phi(x,y)\in\mathbb{R}^{d}$ is a known feature map and $\theta_{t}$ drifts with a total variation budget $V_{T}$, where $V_{T}>0$ is a known constant to the agent.
$\sum^{T}_{t=2}\|\theta_{t}-\theta_{t-1}\|_{2}\leq V_{T}$ (8)
Define the value of the policy at time $t$ as
$J_{t}(\pi):=\mathbb{E}_{y_{t}\sim\pi(\cdot\mid x_{t})}[u_{t}(x_{t},y_{t})].$ (9)
Although EvoDPO is implemented in phases, the same process can be indexed per step $t$ by viewing the reference as piecewise constant over time. We define $\pi^{*}_{t}$ as the instantaneous optimal policy, $\pi_{t}$ as the policy produced by the algorithm at time $t$. We also define the KL-regularized comparator $\pi^{kl}_{t}$ for the reference update.
$\pi_{t}:=\arg\min_{\pi\in\Pi}L_{EvoDPO}(\pi;\pi_{ref,t-1},\mathcal{D}_{t})$ (10)
$\pi^{*}_{t}:=\arg\max_{\pi\in\Pi}J_{t}(\pi)$ (11)
$\pi^{kl}_{t}:=\arg\max_{\pi\in\Pi}\bigl\{J_{t}(\pi)-\beta_{ref}D_{KL}(\pi(\cdot|x_{t})||\pi_{ref,t-1}(\cdot|x_{t}))\bigr\}$ (12)
Dynamic regret decomposition
Under a non-stationary condition, we analyze the decomposed dynamic regret
$r_{t}=\{\underbrace{J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})}_{\text{Reference-induced bias}}\}+\{\underbrace{J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})}_{\text{Learning Error}}\}$ (13)
We also define cumulative dynamic regret as
$R_{T}=\sum^{T}_{t=1}r_{t}.$ (14)
For this theoretical work, we set some assumptions as below:
Assumption 1 (Boundedness)
We assume the feature map, parameter, and utilities are uniformly bounded.
$\|\phi(x,y)\|_{2}\leq\phi_{max},\quad\|\theta_{t}\|_{2}\leq\theta_{max},\quad|u_{t}(x,y)|\leq u_{max}$
Assumption 2 (Reference support)
We assume the reference policy has full support, uniformly lower-bounded by $\pi_{min}$.
$\pi_{ref,t}(y|x)\geq\pi_{min}>0$
Assumption 3 (Oracle switching budget)
Let $B_{T}$ be the total number of oracle-optimal action switches up to time $T$.
$B_{T}\leq\frac{2\phi_{max}}{\gamma}V_{T}$
under a margin condition. Detailed derivation is in Lemma A.3 in Appendix A.
Learning Error.
From the dynamic regret decomposition in (13), the cumulative learning error $R^{error}_{T}$ is defined as:
$R^{error}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})\bigr).$
By Lemma A.6 and Appendix A.4, with probability at least $1-\delta$
$R^{error}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa})$ (15)
where $\kappa\in(0,1)$.
Reference-induced bias.
Define the reference-induced bias as
$R^{bias}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr).$
where $\pi_{t}^{kl}$ is the KL-regularized comparator. As shown in Appendix A.4, this cumulative bias is bounded by the total-variation budget:
$R_{T}^{bias}\leq\mathcal{O}(V_{T}).$
Thus, the cumulative dynamic regret decomposes as $R_{T}=R_{T}^{\mathrm{error}}+R_{T}^{\mathrm{bias}}$. Combining the bias bound with Eq. (15), with probability at least $1-\delta$
$R_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa}+V_{T}).$ (16)
3 Numerical Results
In this section, we conduct a comprehensive set of experiments to evaluate our proposed methods. The main finding is that structured collaboration between a self-evolving research agent and task-distributed multi-LLM supporters enables the ATLAS to solve complex optimization problems that require adaptation to non-stationary environments and time-dependent PDE across varying Reynold number.
(a) Bandit Performance
(b) PINN Performance
Figure 2: Experimental Results across distinct domains. (a) Bandit Negative Mean Regret (NMR). (b) PINN Validation Loss (Log Scale). Shaded regions represent the Standard Error of the Mean (SEM) across 5 independent seeds.
Table 2: Quantitative Summary. We compare the initial and final performance of ATLAS against EvoDPO (ablation study) and EvoTune (baseline). Left column: Negative Mean Regret for the bandit task (higher is better), showing percentage improvement. Right column: Validation Loss for the PINN task (lower is better), showing the fold reduction in loss magnitude.
Bandit PINN
Method Initial Final Improvement Method Initial Final Reduction
EvoTune -0.621 -0.509 18.0% EvoTune 0.880 $1.84\times 10^{-2}$ $48\times$
EvoDPO -0.621 -0.507 18.3% EvoDPO 0.880 $9.76\times 10^{-4}$ $902\times$
[…]
3.1 Setup
Tasks and formulations.
We use two distinct domains to evaluate the framework. Detailed problem formulations and mathematical definitions are provided in the Appendix B.2.
Decision Making.
Non-stationary $k$-armed contextual bandit represents a decision-making problem under concept drift, where the decision-maker must choose the optimal arm while the underlying model parameter changes over time. The decision-maker observes context $x_{t}\in\mathbb{R}^{d}$ and selects one of $k$ arms. The reward depends on a drifting model parameter $\theta_{t}$ and the corresponding action. To rigorously test adaptation, we define the parameter evolution on the unit sphere with a total variation budget $V_{T}$.
$\theta_{t+1}=\frac{\theta_{t}+\delta_{t}}{\|\theta_{t}+\delta_{t}||_{2}}$ (17)
The research agent design strategy for the decision-maker to solve the problem and the supporter agents propose various strategies (e.g., Sliding-Window LinUCB) and dynamically tune hyperparameters (window size $W$ and regularization $\lambda$).
Scientific Machine Learning (SciML).
We consider a partial differential equation (PDE) setting for continuous optimization of PINN loss design, where the agent designs a loss function to solve the 1D viscous Burgers’ equation:
$u_{t}+uu_{x}-\nu u_{xx}=0,\quad x\in[-1,1],t\in[0,1]$ (18)
Instead of relying on fixed, hand-tuned loss weights, our agent explores and optimizes its policy to adaptively reweight the composite PINN loss
$\mathcal{L}(\theta)=\lambda_{pde}\mathcal{L}_{pde}+\lambda_{ic}\mathcal{L}_{ic}+\lambda_{bc}\mathcal{L}_{bc}$ (19)
that adaptively balances these terms during training. To rigorously test adaptation, we introduce a time-varying (dependent) PDE parameter where the viscosity parameter $\nu$ drifts over time, forcing the agent to handle varying degrees of stiffness and shock formation.
Agents.
ATLAS employs a hierarchical multi-agent structure, utilizing four specialized LLMs to drive the self-evolution process (agent roles and prompts are detailed in Appendix B.3 and C)
(i) The Research Agent functions as the primary researcher. Initialized with domain-specific prompts, it iteratively generates a Python solution code.
(ii) The Exploration Supporter acts as a senior technical advisor. It performs static analysis on the research agent’s code to propose architectural patches. It identified and resolved domain-specific bottlenecks, such as reducing the fixed sliding window (e.g., $50\to 20$) for the bandit task and recommending Huber loss with dynamic normalization to stabilize exploding gradients in PINN.
(iii) The Fine-Tuning Strategist regulates the learning dynamics. By monitoring training diagnostics, it dynamically tunes the fine-tuning hyperparameters.
(iv) The Policy Inspector serves as a safety gate for reference updates. It enforces a strict trust region, permitting updates only when the candidate $\pi^{kl}_{k}$ demonstrates sufficient utility improvement while satisfying a trust-region KL divergence constraint.
Baselines.
We evaluate the performance of our full framework, ATLAS, through an ablation study and a comparison against a standard baseline (detailed in Appendix B.1) to isolate the contributions of our proposed components:
(i) EvoDPO: An ablation study of our proposed method that retains adaptivity via adaptive reference updates, but operates as a single agent without the multi-agent architectural benefits.
(ii) EvoTune: A baseline method that utilizes a fixed reference policy and maintains static hyperparameters over the entire course of training.
3.2 Results
Experimental setup.
Our framework is evaluated on both the non-stationary $k$-armed contextual bandit and PINN tasks. All experiments were conducted on a cluster using NVIDIA H100 GPUs; see Appendix B.5 for hyperparameters and hardware details.
Bandit.
For the bandit task, the performance is measured by the Negative Mean Regret (NMR) over the horizon (higher is better). As shown in Figure 2(a) and Table 2 (left), ATLAS achieves the highest score, significantly outperforming both the static baseline (EvoTune) and the ablation without supporters (EvoDPO). Specifically, ATLAS improves the mean initial score from 0.621-0.621 to a final mean of 0.493-0.493 ($20.6\%$ improvement). In comparison, EvoDPO offers a slight advantage over the baseline (converging to 0.507-0.507 vs 0.509-0.509), but eventually plateaus. This performance gap highlights that, although the adaptive update rule of EvoDPO yields measurable improvements, the feedback provided by the Exploration Supporter and the Fine-Tuning Strategist is crucial for systematically adjusting sliding-window parameters and DPO hyperparameters to navigate concept drift.
PINN.
Both Figure 2(b) and Table 2 (right) illustrate the validation loss trajectories and quantitative summary for the PINN task. The objective is to minimize the score (validation loss), defined as the composite loss on the held-out collocation points. ATLAS significantly outperforms the baseline, achieving a final mean validation loss of $3.00\times 10^{-5}$, representing a massive $29,344\times$ improvement over the initial mean score. While EvoDPO significantly improves upon the baseline ($9.76\times 10^{-4}$ vs $1.84\times 10^{-2}$), it exhibits higher variance than the full method. This indicates that while the adaptive update rule effectively navigates the early landscape, the multi-supporter agents layer is critical for consistently overcoming optimization stiffness to achieve high-precision convergence.
Table 3: Activity of the Fine-Tuning Strategist and Policy Inspector. Values represent the mean intervention rate across all trials.
Strategist Intervention Rate Inspector
Task $\Delta\beta$ $\Delta$ Threshold $\Delta$ Epoch Accept Rate
Bandit 22.7% 31.1% 7.3% 16.4%
PINN 16.0% 24.0% 2.0% 38.0%
Agent Dynamics Analysis.
Table 3 summarizes the intervention frequencies of the auxiliary agents. The Fine-Tuning Strategist actively modulated the learning dynamics by intervening on hyperparameters (specifically $\beta$ and scoring thresholds) in roughly 20–30% of the rounds. Notably, the Policy Inspector exhibited distinct behaviors across domains. In the PINN task, the acceptance rate was relatively high ($38.0\%$), reflecting the deterministic nature of the physics gradients, which allowed for consistent policy improvements. Conversely, in the non-stationary bandit task, the acceptance rate was lower ($16.4\%$). This indicates that the Policy Inspector acted as a strict conservatism gate, frequently rejecting updates when the candidate policy would cause overly aggressive updates. Full trial-by-trial dynamics are detailed in Appendix B.4.
3.3 Ablation Study
To isolate the impact of our contributions, we evaluate the system components incrementally in Table 4. The transition from EvoTune to EvoDPO highlights the benefit of adaptivity, which reduces the PINN validation loss by nearly two orders of magnitude ($1.84\times 10^{-2}\to 9.76\times 10^{-4}$). However, the addition of the multi-LLM supporters layer (ATLAS) provides a solution for overcoming stagnation: it enables dynamic parameter adaptation to navigate concept drift in the Bandit task (improving NMR to -0.493) and refines the loss of PINN task by an additional order of magnitude to $3.00\times 10^{-5}$, confirming the synergistic benefit of combining the strategic guidance from auxiliary agents with EvoDPO’s adaptivity.
Table 4: Ablation Study. Comparison of final performance across Bandit and PINN tasks. Bandit values represent Negative Mean Regret (NMR), where higher is better. PINN values represent Validation Loss on held-out collocation points, where lower is better. Metrics are averaged across 5 independent seeds.
Configuration Bandit PINN
Baseline (EvoTune) 0.509-0.509 $1.84\times 10^{-2}$
+ Adaptivity (EvoDPO) 0.507-0.507 $9.76\times 10^{-4}$
+ LLM Supporters (ATLAS) $\mathbf{-0.493}$ $\mathbf{3.00\times 10^{-5}}$
4 Conclusion
We introduced ATLAS, a framework for self-evolving research agents that decomposes the long-horizon evolutionary procedure into role-specialized supporter agents for exploration, strategy control, and policy inspection.
Our primary algorithmic contribution, EvoDPO, performs progressive preference optimization via adaptive reference updates. We provide theoretical guarantees via a non-stationary decision-making problem. On the implementation level, ATLAS demonstrates significant performance, robustness, and reliability when applied to challenging domains such as non-stationary contextual bandits and scientific machine learning.
Future work will extend this framework to collaborative multi-agentic evolution and to broader scientific discovery tasks in which evaluation signals are sparse or noisy.
Impact Statement
Our work on ATLAS and the EvoDPO algorithm provides a pathway to develop self-evolving AI agents capable of tackling complex, non-stationary problems in scientific computing and automated decision-making. By demonstrating how task-distributed multi-agent support can stabilize long-horizon preference learning, we hope to accelerate scientific discovery and automate robust code generation in engineering domains, such as solving partial differential equations or optimizing online learning strategies, where manual tuning is often a bottleneck.
However, as with any framework that enables autonomous self-improvement, there are inherent risks related to objective misalignment and safety. If the reference management mechanisms, specifically the Policy Inspector and trust region constraints, are not rigorously tuned, there is a risk that agents could exploit proxy metrics or manipulate the reward system, leading to solutions that are mathematically high-scoring but physically invalid or unsafe. Additionally, the capability to automate research workflows raises dual-use concerns; the same optimization power used to refine PINNs could potentially be applied to optimize harmful objectives if deployed without safeguards.
Researchers and practitioners utilizing ATLAS should therefore be strictly mindful of the boundaries defined in the exploration and inspection phases, ensuring that physical constraints and ethical guidelines are integrated directly into the supporter agents’ prompts and evaluation logic. Ultimately, by providing a structured, stable, and transparent framework for agentic evolution, we believe this work advances the responsible automation of scientific research and fosters the development of reliable AI systems for complex problem-solving.
Appendix A Proofs for Section 2.3
This appendix provides detailed derivations for the dynamic regret bound of evolving direct preference optimization (EvoDPO).
A.1 Scope and Limitation of the Theoretical Abstraction
Our dynamic regret analysis is intentionally stylized: it studies a non-stationary linear preference contextual bandit under a Bradley–Terry model with a finite action set and linear utilities. This abstraction isolates one mechanism that is common to EvoDPO and long-horizon preference optimization—reference progression under drift—but it does not attempt to faithfully model full sequence-level LLM fine-tuning dynamics. In particular, the analysis does not capture (i) the accept/reject nature of the inspector gate, (ii) restriction to finite candidate sets $C_{k}$ of checkpoints, or (iii) the token-level KL proxy used in practice. These components are treated as algorithmic approximations and are validated empirically via acceptance-rate statistics and sensitivity/robustness diagnostics (App. B.6B.7).
Interpretation.
The theorem should be read as evidence that systematic reference refresh can yield sublinear dynamic regret in a drifting preference environment, whereas a fixed reference can incur linear bias under drift; the gated implementation aims to approximate systematic refresh while enforcing conservative stability constraints, whose reliability we quantify empirically (App. B.6B.7).
A.2 Preliminaries and Definitions
We first formalize the sliding-window estimator and its requisite assumptions. For brevity, write $\phi_{\tau}:=\phi(x_{\tau},y_{\tau})$. Let $\mathcal{W}_{t}$ denote the sliding window of recent history with size $W=|\mathcal{W}_{t}|=T^{\kappa}$ where $\kappa\in(0,1)$ and controllable parameter by decision-maker. We define the regularized covariance matrix $A_{t}$ as:
$A_{t}=\sum_{\tau\in\mathcal{W}_{t}}\phi_{\tau}\phi^{\top}_{\tau}+\lambda I_{d}$ (20)
Estimator
Let $p_{\tau}\in\{0,1\}$ denote the observed preference label at $\tau$. We use the regularized logistic regression estimator:
$\hat{\theta}_{t}=\arg\min_{\theta_{t}}\sum_{\tau\in\mathcal{W}_{t}}\ell\!\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)+\frac{\lambda}{2}\|\theta_{t}\|_{2}^{2}$ (21)
where
$\ell\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)=\log\left(1+e^{\theta_{t}^{\top}\phi_{\tau}}\right)-p_{\tau}\,\theta_{t}^{\top}\phi_{\tau}$ (22)
Oracle-optimal action
Define the oracle-optimal action
$y^{*}_{t}(x)\in\arg\max_{y\in\mathcal{Y}}u_{t}(x,y)$
Assumption A.1 (Covariance Diversity).
There exists a constant $c>0$ such that for all $t\geq W$,
$\lambda_{\min}(A_{t})\geq c\cdot W$
Assumption A.2.
There exists $m_{0}>0$ such that for all relevant $z$,
$\sigma(z)(1-\sigma(z))\geq m_{0}$
We also use that $\sigma$ is $L_{\sigma}$-Lipschitz with $L_{\sigma}\leq\frac{1}{4}$:
$|\sigma(a)-\sigma(b)|\leq L_{\sigma}|a-b|$
A.3 Auxiliary Lemmas
Lemma A.3 (Oracle Switching Budget).
Under the margin condition with gap $\gamma>0$, the number of switches of the oracle-optimal action satisfies
$B_{T}\leq\frac{2\phi_{max}}{\gamma}\sum^{T}_{t=2}\|\theta_{t}-\theta_{t-1}\|_{2}\leq\frac{2\phi_{max}}{\gamma}V_{T}.$
Proof.
Fix $t\geq 2$ and abbreviate $y_{t}^{*}:=y_{t}^{*}(x_{t})$ and $y_{t-1}^{*}:=y_{t-1}^{*}(x_{t})$. This oracle-optimal action is defined in A.2. If $y_{t}^{*}=y_{t-1}^{*}$, there is nothing to show. Suppose $y_{t}^{*}\neq y_{t-1}^{*}$. By optimality at time $t$,
$u_{t}(x_{t},y_{t}^{*})\geq u_{t}(x_{t},y_{t-1}^{*}).$
Using linear utilities and $\|\phi(x,y)\|_{2}\leq\phi_{\max}$, for any $y$,
$|u_{t}(x_{t},y)-u_{t-1}(x_{t},y)|=|\langle\theta_{t}-\theta_{t-1},\phi(x_{t},y)\rangle|\leq\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}.$
Therefore,
$u_{t}(x_{t},y_{t}^{*})\leq u_{t-1}(x_{t},y_{t}^{*})+\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2},$
$u_{t}(x_{t},y_{t-1}^{*})\geq u_{t-1}(x_{t},y_{t-1}^{*})-\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}.$
Combining with $u_{t}(x_{t},y_{t}^{*})\geq u_{t}(x_{t},y_{t-1}^{*})$ gives
$u_{t-1}(x_{t},y_{t}^{*})-u_{t-1}(x_{t},y_{t-1}^{*})\geq-2\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}.$
But by the margin condition at time t1t-1 (for context $x_{t}$), since $y_{t-1}^{*}$ is optimal under $u_{t-1}(x_{t},\cdot)$ and $y_{t}^{*}\neq y_{t-1}^{*}$,
$u_{t-1}(x_{t},y_{t-1}^{*})-u_{t-1}(x_{t},y_{t}^{*})\geq\gamma,$
equivalently $u_{t-1}(x_{t},y_{t}^{*})-u_{t-1}(x_{t},y_{t-1}^{*})\leq-\gamma$. Thus, $\gamma\leq 2\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}$
$\mathbf{1}\{y_{t}^{*}\neq y_{t-1}^{*}\}\leq\frac{2\phi_{\max}}{\gamma}\|\theta_{t}-\theta_{t-1}\|_{2}.$
Summing over $t=2,\dots,T$ yields the claim. ∎
Lemma A.4 (Local variation bound).
Define the local window variation
$V_{t,W}:=\sum^{t-1}_{j=t-W}\|\theta_{j+1}-\theta_{j}\|_{2}.$
Then,
$\sum_{\tau\in\mathcal{W}_{t}}\|\theta_{t}-\theta_{\tau}\|_{2}\leq WV_{t,W},\quad\sum^{T}_{t=1}V_{t,W}\leq WV_{T}.$
Proof.
For any $\tau\in\mathcal{W}_{t}$, $\|\theta_{t}-\theta_{\tau}\|_{2}\leq V_{t,W}$. Therefore, $\sum_{\tau\in\mathcal{W}_{t}}\|\theta_{t}-\theta_{\tau}\|_{2}\leq WV_{t,W}$.
For the second claim, note that each increment $\|\theta_{j+1}-\theta_{j}\|_{2}$ appears in at most $W$ windows $\{V_{t,W}\}_{t=1}^{T}$. Therefore,
$\sum_{t=1}^{T}V_{t,W}=\sum_{t=1}^{T}\sum_{j=t-W}^{t-1}\|\theta_{j+1}-\theta_{j}\|_{2}\leq\sum_{j=1}^{T-1}W\,\|\theta_{j+1}-\theta_{j}\|_{2}=WV_{T}.$
Lemma A.5 (Self-normalized inequality).
Based on Assumption 1, $\|\phi_{\tau}\|_{2}\leq\phi_{max}$ and $\eta_{\tau}$ are zero-mean 1-sub-Gaussian. Then with probability at least $1-\delta$,
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda+W\phi_{\max}^{2}}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}.$
Proof.
We use self-normalized inequality (Lemma 9) derivation from (Abbasi-yadkori et al., 2011).
$\log\!\left(\frac{\det(A_{t})^{1/2}\det(\lambda I)^{-1/2}}{\delta}\right)=\frac{1}{2}\log\!\left(\frac{\det(A_{t})}{\det(\lambda I)}\right)+\log\!\left(\frac{1}{\delta}\right).$
Using the determinant bound,
$\det(A_{t})\leq\left(\lambda+\frac{W\phi_{\max}^{2}}{d}\right)^{d},\qquad\det(\lambda I)=\lambda^{d},$
hence
$\frac{1}{2}\log\!\left(\frac{\det(A_{t})}{\det(\lambda I)}\right)\leq\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right).$
By combining determinant-trace bound condition, we get follwoing inequality by self-normalized vector martingale with probability at least $1-\delta$
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{A_{t}^{-1}}\leq\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}.$
Convert to Euclidean norm:
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda_{\max}(A_{t})}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}$
Substitute $\lambda_{\max}(A_{t})\leq\lambda+W\phi_{\max}^{2}$:
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda+W\phi_{\max}^{2}}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\quad w.p\quad 1-\delta.$
Lemma A.6 (Parameter estimation error bound).
We use Assumption 1, A.1, A.2, Let $\eta_{\tau}=p_{\tau}-\sigma\bigl(\langle\theta_{\tau},\phi_{\tau}\rangle\bigr)$ be zero-mean 1-sub-Gaussian noise. Then with probability at least $1-\delta$,
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\left[\frac{\phi_{\max}^{2}L_{\sigma}V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\right]$
Proof.
Define the objective
$L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\ell\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)+\frac{\lambda}{2}\|\theta_{t}\|_{2}^{2}$
We use the chain rule on $L_{t}(\theta_{t})$ to get the gradient and Hessian. Then
$\nabla L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\bigl(\sigma(\langle\theta_{t},\phi_{\tau}\rangle)-p_{\tau}\bigr)\,\phi_{\tau}+\lambda\theta_{t}.$
$\nabla^{2}L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\sigma(\langle\theta_{t},\phi_{\tau}\rangle)\bigl(1-\sigma(\langle\theta_{t},\phi_{\tau}\rangle)\bigr)\,\phi_{\tau}\phi_{\tau}^{\top}+\lambda I$
By Assumptions A.1, A.2, $\lambda_{min}(\nabla^{2}L_{t}(\theta_{t}))\geq m_{0}cW$. Using strong convexity
$\big\langle\nabla L_{t}(\hat{\theta}_{t})-\nabla L_{t}(\theta_{t}),\;\hat{\theta}_{t}-\theta_{t}\big\rangle\geq\lambda_{\min}\!\big(\nabla^{2}L_{t}(\theta_{t})\big)\cdot\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}\geq m_{0}cW\,\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}.$
Since the optimal $\hat{\theta}_{t}$ is computed by $\nabla_{\theta}L_{t}(\hat{\theta}_{t})=0$, we substitute this optimality to strong convexity inequality above. Then, we get following inequality
$m_{0}cW\,\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}\;\leq\;\big\langle-\nabla L_{t}(\theta_{t}),\;\hat{\theta}_{t}-\theta_{t}\big\rangle\;\leq\;\|\nabla L_{t}(\theta_{t})\|_{2}\cdot\|\hat{\theta}_{t}-\theta_{t}\|_{2}.$
Therefore, the upper bound of estimate error is:
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\frac{1}{m_{0}cW}\,\|\nabla L_{t}(\theta_{t})\|_{2}.$
Rewrite using drift and noise:
$\nabla L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}-\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}+\lambda\theta_{t},$
Take norms and apply triangle inequality:
$\|\nabla L_{t}(\theta_{t})\|_{2}\leq\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}\Bigr\|_{2}+\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}+\lambda\|\theta_{t}\|_{2}.$
Since $\sigma(\cdot)$ is $L_{\sigma}$-Lipschitz, the drift term satisfies
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}\Bigr\|_{2}\leq\sum_{\tau\in\mathcal{W}_{t}}L_{\sigma}\,\|\theta_{t}-\theta_{\tau}\|_{2}\,\|\phi_{\max}\|_{2}^{2}\leq L_{\sigma}\phi^{2}_{max}W\;V_{t,W}$
Now, we bound the second term via the (Li et al., 2010) log-determinant argument. As we defined the $A_{t}$ matrix in (16),
$\lambda_{\max}(A_{t})\leq\lambda+\lambda_{\max}\!\left(\sum_{\tau\in\mathcal{W}_{t}}\phi_{\tau}\phi_{\tau}^{\top}\right)\leq\lambda+\sum_{\tau\in\mathcal{W}_{t}}\|\phi_{\tau}\|_{2}^{2}\leq\lambda+W\phi_{\max}^{2}.$
The noise term is bounded by Lemma A.5, and $\lambda\|\theta_{t}\|_{2}\leq\lambda\theta_{\max}$.
$\|\nabla L_{t}(\theta_{t})\|_{2}\leq\left[{L_{\sigma}\phi_{\max}^{2}\,W\,V_{t,W}}\;+\;\sqrt{\lambda+W\phi_{\max}^{2}}\,\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\lambda\theta_{\max}\right]$
Dividing by $m_{0}cW$ yields the result.
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\left[\frac{L_{\sigma}\phi_{\max}^{2}\,V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\right].$
Lemma A.7 (KL divergence bound).
Let $u_{t}(x_{t},y_{t})=\langle\theta_{t},\phi(x_{t},y_{t})\rangle$ and $\hat{u}_{t}(x_{t},y_{t})=\langle\hat{\theta}_{t},\phi(x_{t},y_{t})\rangle$. The KL-divergence between KL-regularized policy $\pi^{kl}_{t}$ and the agent policy $\pi_{t}$ is bounded by parameter estimate error:
$D_{KL}(\pi^{kl}_{t}(\cdot|x_{t})||\pi_{t}(\cdot|x_{t}))\leq\frac{\phi_{max}^{2}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}^{2}}{2\beta^{2}}.$
Proof.
From the definition of KL-divergence:
$D_{KL}\!\big(\pi^{kl}_{t}(\cdot|x_{t})\,\|\,\pi_{t}(\cdot|x_{t})\big)=\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\,\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}.$
We use the policy definition, which is defined in (Rafailov et al., 2024)
$\pi^{kl}_{t}(y_{t}|x_{t})=\frac{1}{Z(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}$
$Z(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}u_{t}(x_{t},y_{t}))$
$\pi_{t}(y_{t}|x_{t})=\frac{1}{\hat{Z}(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t})}$
$\hat{Z}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t}))$
From the KL-divergence definition above, we derive the log term first.
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\frac{\frac{1}{Z(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}}{\frac{1}{\hat{Z}(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t})}}$
$\log\frac{\hat{Z}(x_{t})}{Z(x_{t})}+\frac{1}{\beta}(u_{t}(x_{t},y_{t})-\hat{u}_{t}(x_{t},y_{t}))$
Let $u_{t}(x_{t},y_{t})-\hat{u}_{t}(x_{t},y_{t})=\Delta u_{t}$. Then,
$\hat{Z}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}(u_{t}(x_{t},y_{t})-\Delta u_{t}))$
$Z(x_{t})\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}(\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t})))$
$\frac{\hat{Z}(x_{t})}{Z(x_{t})}=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t})}{\beta}}$
Thus, the log term can be concluded as below
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\mathbb{E}_{y_{t}\sim\pi^{kl}(\cdot|x_{t})}[\exp{(-\frac{\Delta u_{t}}{\beta})}]+\frac{1}{\beta}(\Delta u_{t})$
Now we take the KL-divergence definition again
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}\biggr]=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\exp{\bigl(-\frac{\Delta u_{t}}{\beta}\bigr)}\biggr]+\frac{1}{\beta}\mathbb{E}_{\pi^{kl}}[\Delta u_{t}]$
Let $\frac{\Delta u_{t}}{\beta}=X$. Then,
$D_{KL}=-\mathbb{E}[X]+\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[e^{-(X-\mathbb{E}[X])}\biggr]+\mathbb{E}[X]$
Thus, $D_{KL}$ can be simplified as below
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\bigl[e^{-(X-\mathbb{E}[X])}\bigr]$
Now, we derive the upper bound of $\Delta u_{t}$ with parameter estimate error. Since we defined $u_{t}$ in (3),
$\Delta u_{t}(y)=\theta^{\top}_{t}\phi(x_{t},y_{t})-\hat{\theta}^{\top}_{t}\phi(x_{t},y_{t})\leq\phi_{max}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}$
With the above inequality, we can define the range
$\frac{\Delta u_{t}}{\beta}\in\biggl[-\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta},\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta}\biggr]$
Let $a:=\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta}$ so that $X\in[-a,a]$. Since $\mathbb{E}[X]\in[-a,a]$, the centered variable satisfies
$\max(X-\mathbb{E}[X])-\min(X-\mathbb{E}[X])\leq 2a.$
By Hoeffding’s lemma, with s=1s=-1,
$\log\mathbb{E}\big[e^{-(X-\mathbb{E}[X])}\big]\leq\frac{(2a)^{2}}{8}=\frac{a^{2}}{2}.$
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))\leq\frac{\phi_{max}^{2}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}^{2}}{2\beta^{2}}$
A.4 Proof of regret analysis
We decompose the cumulative dynamic regret $R_{T}$ into Learning Error ($R^{error}_{T}$) and Reference Bias ($R^{bias}_{T}$).
Learning error
Now, we derive $J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})$.
$r_{t}=J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})=\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})$
$\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})\leq\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|$
$\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot|u_{t}(x_{t},y_{t})|$
Since $\|u_{t}(x_{t},\cdot)\|_{\infty}:=\max_{y_{t}\in\mathcal{Y}}|u_{t}(x_{t},y_{t})|$, we get the following inequality
$\leq\|u_{t}(x_{t},\cdot)\|_{\infty}\sum_{y_{t}\in\mathcal{Y}}\Bigl|\pi_{t}^{kl}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\Bigr|$
$\text{norm inequality}\quad\|u\|_{\infty}\leq\|u\|_{2}\leq\|u\|_{1}$
We can write again as
$r_{t}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{\infty}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{2}$
We convert $\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|$ into L1 norm $\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}$ and using Assumption 1, derive the upper bound as below
$r_{t}\leq\theta_{max}\cdot\phi_{max}\cdot\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}$
Since the $\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})$ is measured in L1 norm, we convert the term into a KL-divergence term using Pinsker’s inequality and Lemma A.7,
$J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})\leq\sqrt{2D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))}\leq\frac{\theta_{max}\phi^{2}_{max}}{\beta}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}$
Summing over $t$ and applying Lemma A.6, together with $\sum^{T}_{t=1}V_{t,W}\leq W\;V_{T}$,
$R^{error}_{T}\leq\sum^{T}_{t=1}\frac{\theta_{max}\phi^{2}_{max}}{\beta}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}.$
From Lemma A.6, we finalize the cumulative learning error bound
$R^{error}_{T}\leq\frac{T\;\theta_{max}\phi^{2}_{max}}{\beta}\biggl[\frac{L_{\sigma}\phi_{\max}^{2}\,\,V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\biggr]$
Since total variation is bounded by $V_{T}$ and $W=T^{\kappa}$, we write this upper bound with $\mathcal{O}$ function dependent on $T$ term, with probability at least $1-\delta$
$R^{error}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa}).$
Reference Bias
We take the definition of $\pi^{kl}(y_{t}|x_{t})$ again
$\pi^{kl}_{t}(y_{t}|x_{t})=\frac{1}{Z(x)}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}$
$\log\pi^{kl}_{t}(y_{t}|x_{t})=\log\pi_{ref,t}(y_{t}|x_{t})+\frac{1}{\beta}u_{t}(x_{t},y_{t})-\log Z(x_{t})$
$\frac{1}{\beta}u_{t}(x_{t},y_{t})=\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t}).$
Since we defined $J_{t}(\pi)$ function,
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\frac{u_{t}(x_{t},y_{t})}{\beta}$
substitute above $\frac{1}{\beta}u_{t}(x_{t},y_{t})$, then we get
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr].$
To write the above equation in terms of KL-divergence, we decompose the log term as below
$\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}=\log\frac{\pi(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}-\log\frac{\pi(y_{t}|x_{t})}{\pi^{kl}_{t}(y_{t}|x_{t})}.$
Substitute this decomposed term to $\frac{1}{\beta}J_{t}(\pi)$.
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}-\log\frac{\pi(y_{t}|x_{t})}{\pi^{kl}_{t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr]$
$\frac{1}{\beta}J_{t}(\pi)=D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t}).$
Then, $J_{t}(\pi)$ is redefined as below
$J_{t}(\pi)=\beta\bigl(D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t})\bigr).$
Now, we substitute $\pi=\pi^{kl}_{t}$ and $\pi=\pi^{*}_{t}$. Then,
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})=\beta\bigl(D_{KL}(\pi^{*}_{t}||\pi_{ref,t})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})-D_{KL}(\pi^{kl}_{t}||\pi_{ref,t})\bigr)$
Our framework uses EvoDPO, which updates the current reference policy to the previous KL-regularized optimal policy (only for EvoDPO; if it is DPO, $\pi_{ref,t}=\pi_{ref}$). Thus, $\pi_{ref,t}=\pi^{kl}_{t-1}$. Since the third term $D_{KL}(\pi^{kl}_{t}||\pi^{kl}_{t-1})$ is a non-negative term, this term is dropped out for the upper bound.
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\leq\beta\bigl(D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})\bigr)$
For cumulative regret analysis,
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})-\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{1})\dots+\beta D_{KL}(\pi^{*}_{T}\|\pi^{kl}_{T-1})-\beta D_{KL}(\pi^{*}_{T}\|\pi^{kl}_{T})$
Thus, this term is simplified again
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]$
We analyze the second term of the RHS. Since the deterministic policy is $\pi^{*}$ that selects action $y^{*}$, $D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})$ is simplified as below
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})=\sum_{y_{t}}\pi^{*}_{t}(y^{*}_{t}|x_{t})\log\frac{\pi^{*}_{t}(y^{*}_{t}|x_{t})}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}=1\cdot\log\frac{1}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}$
$D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})$ term is also simplified in same way
$D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})=1\cdot\log\frac{1}{\pi^{kl}_{t-1}(y^{*}_{t-1}|x_{t})}$
Then
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})=\log\frac{\pi^{kl}_{t-1}(y^{*}_{t-1}|x_{t})}{\pi^{kl}_{t-1}(y^{*}_{t}|x_{t})}$
We take the definition of each numerator and denominator again
$D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\leq\log\frac{\pi_{ref,t-1}(y^{*}_{t-1}|x_{t})}{\pi_{ref,t-1}(y^{*}_{t}|x_{t})}+\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}$
From assumption 2, whenever a switch happens,
$\log\pi_{ref,{t-1}}(y^{*}_{t-1}|x_{t})-\log\pi_{ref,t-1}(y^{*}_{t}|x_{t})\leq\log 1-\log\pi_{min}\leq\log\frac{1}{\pi_{min}}$
If a switch has not happened, the $\log$ term is exactly 0. Thus, this $\log$ term follows the cumulative reference mismatch term, which is described in assumption 3; the total number of oracle-optimal action switch bound condition
$\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]\leq B_{T}\cdot\log\frac{1}{\pi_{min}}$
Since the second term $\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}$ is environment drift term and cumulated drift is bounded by $V_{T}$,
$\sum^{T}_{t=1}\frac{\phi_{max}}{\beta}\|\theta_{t}-\theta_{t-1}\|_{2}\leq\frac{\phi_{max}}{\beta}V_{T}$
From the Lemma A.3, we derived the upper bound of $B_{T}$
$B_{T}\leq\frac{2\phi_{max}}{\gamma}V_{T}$
Since the cumulative reference mismatch bias term in total time $T$ is
$\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\sum^{T}_{2}\bigl[D_{KL}(\pi^{*}_{t}\|\pi^{kl}_{t-1})-D_{KL}(\pi^{*}_{t-1}\|\pi^{kl}_{t-1})\bigr]$
$\leq\beta D_{KL}(\pi^{*}_{1}\|\pi^{kl}_{0})+\beta\biggl(\frac{2\phi_{max}V_{T}}{\gamma}\cdot\log\frac{1}{\pi_{min}}+\frac{\phi_{max}V_{T}}{\gamma}\biggr)$
Therefore, the cumulative reference mismatch bias term is
$R^{bias}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr)\leq\mathcal{O}(V_{T})$
Final bound.
Combining the learning error with the reference bias we obtain:
$R_{T}=R^{error}_{T}+R^{bias}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa}+V_{T})\quad w.p\quad 1-\delta.$
Appendix B Additional Information of Evaluation Setup for Section 3
B.1 Detailed Method Used Formulations
To validate the effectiveness of ATLAS, we compare our proposed method against a standard single-agent baseline:
EvoTune (Baseline): This configuration represents the standard self-evolving agent without external supporter agent guidance. The EvoTune generates solutions via standard DPO (mode=vanilla) on the self-generated data. Crucially, EVOTUNE_USE_TEACHER=0 is set, meaning the Exploration Supporter is disabled (no static analysis feedback) and the Strategist is disabled (use_dpo_strategy_teacher=0), forcing the agent to rely solely on the scalar scores from the evaluator.
EvoDPO (Ablation): This isolates the impact of our proposed Evolving DPO (EvoDPO) algorithm (mode=adaptive) without the full agentic support of ATLAS. Unlike EvoTune (Baseline), the reference model $\pi_{\text{ref,t}}$ is not static; it updates to track the evolving policy $\pi_{t}$. However, unlike ATLAS, the Policy Inspector is disabled (or set to always accept), and the Supporters are disabled (EVOTUNE_USE_TEACHER=0), meaning the reference model updates without KL trust-region safeguard and qualitative feedback. This serves as an ablation to test whether the ”moving reference” mechanism alone is sufficient without the Policy Inspector gating.
ATLAS (Ours): Our full proposed framework. It utilizes the complete task-distributed multi-agent system (Exploration Supporter, Fine-Tuning Strategist ,and Policy Inspector) and distinguishes itself by replacing standard DPO with our Evolving DPO algorithm (mode=adaptive). This enables the reference model $\pi_{\text{ref,t}}$ to evolve over time. Crucially, every update is gated by the Policy Inspector, which rejects updates that violate the drift constraint ($\widehat{\mathrm{KL}}_{k}\leq\delta_{H}$) or fail to improve validation performance.
B.2 Detailed Problem Formulations and Evaluation Metrics
B.2.1 Non-Stationary Contextual Bandit
We consider a bandit with $k$ arms and context dimension $d$ over a horizon of $H$ steps. At each step $t$, the decision-maker observes context vectors ${(x_{t,a})}_{a=1}^{K}\in\mathbb{R}^{d}$ (normalized to unit norm) and selects an arm $a_{t}$. The reward is $r_{t}=\langle\theta_{t},x_{t,a_{t}}\rangle+\eta_{t}$, with 1-sub gaussian noise $\eta_{t}$. The parameter $\theta_{t}$ drifts with a total variation budget $V_{T}$, evolving as:
[…]
Evaluation. The Research Agent simulates a bandit problem for each candidate policy. The system evaluates performance using the Negative Mean Regret (NMR) over the horizon $H$. This scoring metric aligns with a maximization objective (higher is better):
$\text{score}=-\frac{1}{H}\sum_{t=1}^{H}(r^{*}_{t}-r_{t})$
where $r^{*}_{t}$ is the expected reward of the optimal arm and $r_{t}$ is the expected reward of the selected arm.
B.2.2 1D Burgers’ Equation)
We solve the viscous Burgers’ equation on the domain $(x,t)\in[-1,1]\times[0,1]$:
$u_{t}+uu_{x}-\nu(t)u_{xx}=0$
with initial condition $u(0,x)=-\sin(\pi x)$ and zero Dirichlet boundary conditions u(t,1)=u(t,1)=0u(t,-1)=u(t,1)=0. The Physics-Informed Neural Network (PINN) $u_{\theta}(x,t)$ is trained to minimize a composite loss:
$\mathcal{L}(\theta)=\lambda_{pde}\mathcal{L}_{pde}(\theta)+\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)$
where $\mathcal{L}_{pde}$, $\mathcal{L}_{ic}$, and $\mathcal{L}_{bc}$ are the mean squared errors (MSE) of the PDE residuals, initial condition loss, and boundary condition loss, respectively. The Research Agent’s task is to distribute the weighting coefficients $\lambda_{pde},\lambda_{ic},\lambda_{bc}$ to balance these competing objectives.
Physical Interpretation of Parameter Drift.
A key challenge in PINN training is the dependency on the Reynolds number $Re\propto 1/\nu$. Low Reynolds numbers (high $\nu$) correspond to diffusion-dominated regimes where solutions are smooth and easy to train. High Reynolds numbers (low $\nu$) correspond to convection-dominated regimes where sharp gradients (shocks) form, creating a “stiff” optimization landscape that often leads to failure modes like the trivial zero solution or spectral bias. To evaluate the research agent’s ability to adapt to these changing physical regimes, we implement a drifting viscosity schedule:
$\nu(t)=\nu_{base}\cdot\left(1+\alpha\sin\left(\frac{2\pi t}{T}\right)\right)$
where $\nu_{base}\approx 0.00318$ ($Re\approx 600$), $\alpha=0.9$. This drifts the system between a viscous regime ($\nu\approx 0.006,Re\approx 300$) and a highly turbulent-like shock regime ($\nu\approx 0.0003,Re\approx 6600$). This forces the Research Agent to design a loss strategy that is robust not just to a single static problem, but to the entire transition from smooth flow to shock formation.
Evaluation. We instantiate a fresh neural network (MLP with 4 hidden layers of 32 units) for each candidate loss strategy and train it for 500 optimization steps using the Adam optimizer. We then evaluate the performance on a newly sampled, held-out set of $N_{coll}$ collocation points. We report the Negative Validation Loss (NVL) as the score:
$\text{score}=-\left|\mathcal{L}(\theta;\mathcal{D}_{val})\right|$
where $\mathcal{L}$ is the candidate’s loss function evaluated on the validation set. Scores closer to 0 indicate that the resulting PINN solution satisfies the agent-defined time-dependent physical dynamic setting.
Note: To adhere to standard visualization conventions, the main body of this paper reports the magnitude of this score (i.e., $|\text{score}|$) as “Validation Loss,” where lower is better.
B.3 Detailed Role Definition
B.3.1 Non-Stationary Contextual Bandit
Self-Evolving Research Agent.
The Research Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Research Agent in online learning and contextual bandits. (see the adaptation details in Appendix C.3). For each round, it is provided with the problem context and instructed to implement a Python function policy(context, history, t, ...) that selects an action index (to select arm) $a_{t}\in[0,K-1]$ which will be evaluated by the generated policy and returns the score.
Task-Distributed Multi-LLM Supporters.
The Exploration Supporter (gpt-oss:120b) analyzes the generated policy code. Specifically, the Sliding-Window LinUCB implementation dynamically suggests hyperparameter adjustments. For instance, while the initial window size is set to 50, the supporter may suggest increasing it (e.g./ to 100 or 200) during stable phases to capture longer-term dependencies, or reducing it to adapt more rapidly during high-drift phases. Similarly, it tunes the regularization parameter (e.g., from 1.0 to 0.1) and exploration scalar based on recent performance. In one instance, it flagged a policy with a large fixed window: “The default sliding-window size ($\approx 50$) is too large for a rapidly drifting environment, so the ridge estimator incorporates stale data.” It explicitly patched the code to: if t == 0: policy.window_size = 20 policy.lambda_reg = 0.1 This reduced the estimator’s ”memory” to track the random walk $\theta_{t}$ more effectively.
The Fine-Tuning Strategist (deepseek-r1:32b) monitors training diagnostics during the periodic finetuning updates (every 20 rounds) to adjust DPO hyperparameters (score threshold, $\beta$, and epochs) to prevent the model from overfitting to stale bandit feedback. When it was observed that the student was generating repetitive code (low entropy), it issued a JSON command to set the DPO beta ($\beta$) to $0.1$ and adjust the score_threshold to include a broader range of ”good enough” examples, forcing the model to explore alternative ridge hyperparameters.
The Policy Inspector (llama4:latest) acts as a safety gate for EvoDPO, permitting reference policy updates only when the candidate $\pi^{kl}_{t}$ shows sufficient score improvement ($\Delta_{t}\geq-0.0007$) and satisfies a trust-region KL divergence constraint ($\widehat{\mathrm{KL}}_{k}(\pi^{kl}_{k}||\pi_{\text{ref},k})\leq 0.002$). The logs show it rejecting a candidate update that improved training regret but had a high KL divergence of $0.0045$ on the generated tokens, preventing a ”catastrophic update” where the agent would have dissolved its stable baseline policy in favor of a high-variance experimental one.
B.3.2 Physics-Informed Neural Network (Burgers’ Equation)
Self-Evolving Research Agent.
The Research Agent is initialized with Llama-3.2-1B-Instruct and receives a system prompt defining it as an Research Agent in PDE-constrained optimization. It is tasked with implementing a Python function pinn_loss_strategy(residuals, ic_err, bc_err, extras) that returns a scalar loss tensor. Note that the agent does not solve the optimization problem itself; instead, it defines the scalar objective function that the external optimizer (Adam) minimizes. The prompt explicitly encourages the agent to use adaptive weighting (based on extras[’epoch’]) and to address numerical stiffness (e.g., by re-weighting or normalizing residuals near shocks).
Task-Distributed Multi-LLM Supporters.
The Exploration Supporter (gpt-oss:120b) critiques the generated loss strategy code, identifying issues like scale imbalance between error terms and poor schedule design. For example, it detected that the PDE residual term ($\sim 10^{-2}$) often overshadowed boundary condition errors ($\sim 10^{-5}$), causing ”boundary drift.” In the logs, it explicitly recommended dynamic normalization and component-wise clipping (Huber loss) to fix this: “Replace global clamp with a Huber loss to prevent exploding gradients… Keep a running estimate of the standard deviation of each component and divide by it.” This ensured the optimization respected physical boundaries at $x=\pm 1$.
The Fine-Tuning Strategist (deepseek-r1:32b) dynamically tunes DPO hyperparameters during the periodic fine-tuning updates (every 20 rounds) to ensure the generator learns robust loss structures without collapsing to degenerate solutions. The logs show it actively adjusting $\beta$ (e.g., setting beta=0.7 then 0.5). While it occasionally suggests changes to num_train_epochs or score_threshold (e.g., ”Increasing beta and epochs for better exploration”), the system prioritizes its $\beta$ adjustments to stabilize convergence without inducing mode collapse.
The Policy Inspector (llama4:latest) enforced the improvement threshold. In one instance, it rejected an update where the candidate loss strategy minimized the PDE residual to $10^{-6}$ but caused the Initial Condition error to spike to $10^{-1}$ (a common mode collapse in PINNs known as the ”trivial solution” where $u(t,x)=0$). The inspector flagged this as “Regression: IC error spike” and reverted the reference model, saving the training run from divergence.
B.4 Detailed Agent Dynamics
Table 5: Detailed activity of the Fine-Tuning Strategist and Policy Inspector. Values indicate the fraction of active fine-tuning rounds where a specific intervention was triggered.
Strategist Intervention Rate Inspector
Task Trial $\Delta\beta$ $\Delta$ Threshold $\Delta$ Epoch Acceptance Rate
Bandit 1 4/25 ($16\%$) 10/25 ($40\%$) 2/25 ($8\%$) 1/25 ($4.0\%$)
2 3/22 ($13.6\%$) 9/22 ($40.9\%$) 2/22 ($9.1\%$) 8/22 ($36.4\%$)
3 7/23 ($30.4\%$) 6/23 ($26.1\%$) 1/23 ($4.3\%$) 1/23 ($4.3\%$)
4 6/22 ($27.3\%$) 6/22 ($27.3\%$) 1/22 ($4.5\%$) 7/22 ($31.8\%$)
5 5/19 ($26.3\%$) 4/19 ($21.1\%$) 2/19 ($10.5\%$) 1/19 ($5.3\%$)
Mean 22.7% 31.1% 7.3% 16.4%
PINN 1 0/10 ($0\%$) 0/10 ($0\%$) 0/10 ($0\%$) 5/10 ($50\%$)
2 3/10 ($30\%$) 4/10 ($40\%$) 1/10 ($10\%$) 7/10 ($70\%$)
3 4/10 ($40\%$) 1/10 ($10\%$) 0/10 ($0\%$) 1/10 ($10\%$)
4 0/10 ($0\%$) 3/10 ($30\%$) 0/10 ($0\%$) 1/10 ($10\%$)
5 1/10 ($10\%$) 4/10 ($40\%$) 0/10 ($0\%$) 5/10 ($50\%$)
Mean 16.0% 24.0% 2.0% 38.0%
Table 5 presents the detailed breakdown of Fine-Tuning Strategist interventions and Policy Inspector decisions across 5 independent trials. The variability in the Bandit task highlights the adaptive nature of the ATLAS framework. For example, the number of active fine-tuning rounds varies across trials (e.g., Trial 1 had 25 rounds, while Trial 5 had 19). This variance occurs because ATLAS employs an efficiency mechanism: if the exploration phase yields no improvement in score, the computationally expensive fine-tuning loop is skipped for that round. In contrast, the PINN task consistently achieved score improvements, triggering the fine-tuning loop in all 10 scheduled phases across all trials.
B.5 Hyperparameters, Hardware, and Timing
Hyperparameters. Lists the hyperparameters used for training, and task-specific configurations are summarized in Table 6.
Category Parameter Value
General Total Rounds $501$ (Bandit); $201$ (PINN)
Number of Seeds 55 (Independent Runs)
Sampling Strategy Continuous ($K=20$ rounds)
Number of Islands 66
DPO Training Finetuning Frequency Every $20$ rounds
Initial Learning Rate $5\times 10^{-6}$
Initial Beta ($\beta$) $0.6$
Initial Epochs per Update 11
LoRA Rank / Alpha $r=64,\alpha=32$
KL Thresholds ($\delta_{High}$) $0.002$
Score improvement tolerance ($\epsilon_{s}$) $0.0007$
LLM Temperature Llama-3.2-1B-Instruct 0.9
Gpt-oss: 120b 0.7
Deepseek-r1:32b 0.7
Llama4:latest 0.7
Bandit Task Number of arms ($k$) 5
Dimensions ($d$) 5
Horizon ($H$) 2000
Total variation budget ($V_{T}$) $8000$
Drift limits ($\delta_{min},\delta_{max}$) $1.0,5.0$
PINN Task Initial Viscosity ($\nu$) $0.01/\pi$
Sampling ($N_{coll},N_{ic},N_{bc}$) $512,64,64$
Inner Optimizer Adam
Inner Learning Rate $1\times 10^{-4}$
Inner Optimization Steps $500$
Table 6: Hyperparameters for Training, DPO, and Tasks.
Hardware. For benchmarking each task, we were using 8 Nvidia H100 GPUs with 80 GB of RAM. From this cluster, we specifically allocated our developing research agent (Llama-3.2-1B-Instruct), which used 1 Nvidia H100 GPU with 40 GB of RAM and 14 CPU cores.
Timing. A full run of ATLAS, EvoDPO, and evoTune takes approximately 2 hours per seed for the Bandit task and approximately 1 hour per seed for the PINN task with the provided hardware environment. The runtime for other methods and/or benchmarks falls within a similar range.
B.6 KL Proxy and Limitations
The gate uses a trajectory-conditioned token-level KL proxy (Eq. (2)), computed from next-token distributions under teacher forcing. This proxy is efficient and targets local drift in autoregressive behavior, but it is not equal to unconditional sequence-level KL and can exhibit variance depending on the chosen trajectory and the sampled gate subset $G_{k}$.
B.7 Selection Adaptivity
Because candidates and scores are generated on-policy, gate decisions can be adaptive to the current phase buffer. We mitigate this by evaluating the gate on a fresh randomly sampled subset $G_{k}$ each phase, reducing reuse of the same examples for candidate selection and acceptance decisions.
B.8 Implementation Details
Table 7: Key software libraries and versions.
Category Libraries & Versions
Core Frameworks Python 3.10
PyTorch 2.8.0
CUDA 12.6.0
JAX 0.4.30
LLM & Optimization Transformers 4.57.3
TRL (Transformer Reinforcement Learning) 0.24.0
PEFT 0.17.1
Accelerate 1.10.1
Tokenizers 0.22.1
OpenAI 2.8.1
Scientific Computing NumPy 1.26.4
SciPy 1.13.1
SymPy 1.14.0
Pandas 2.3.3
Our experiments were implemented using PyTorch and Hugging Face Transformers for LLM training. Key libraries and their versions are listed in Table 7.
Appendix C Prompt Templates for Code Generation
This section provides templates for the given prompts for LLM agents. All tasks use the same Fine-Tuning Strategist and Policy Inspector prompt, whereas Exploration Supporter and Self-Evolving Research Agent use suitable prompts for each task.
C.1 Problem Description: Non-stationary Contextual Bandit
Non-stationary linear contextual bandit: K arms, d-dim contexts, drifting $\theta$(t).
[…]
Reward is generated by the evaluator: r_t = context[a_t] * $\theta$(t) + noise.
Goal: minimize regret. You only choose a_t using history (no access to $\theta$ or expected rewards).
The evaluator maintains a sliding-window ridge estimator:
[…]
and exposes it to your policy as:
- policy.A (dxd), policy.b (d,)
Your policy must tune hyperparameters by setting:
- policy.lambda_reg (ridge regularizer, float >> 0)
- policy.window_size (sliding window size W, int >>= 1)
C.2 Problem Description: PINN Burgers
We are evolving loss functions for a physics-informed neural network (PINN) that solves the 1D viscous Burgers equation: u_t + u * u_x - nu * u_xx = 0, (t, x) in [0, 1] x [-1, 1] with:
- Initial condition: u(0, x) = -sin(pi * x)
- Boundary conditions: u(t, -1) = 0, u(t, 1) = 0
The candidate function pinn_loss_strategy(residuals, ic_err, bc_err, extras) receives:
- residuals: squared PDE residuals at N_coll collocation points
- ic_err : squared initial-condition errors at N_ic points
- bc_err : squared boundary-condition errors at 2*N_bc points
- extras : metadata such as epoch and nu
The strategy must combine these into a scalar loss tensor. During evaluation, we compute a scalar total_err = pinn_loss_strategy(...) on held-out samples, then convert it to a scalar score:
err_mag = abs(total_err)
score = -err_mag
So:
- Smaller |total_err| =>> larger score (closer to 0, still negative).
- Larger |total_err| =>> more negative score.
The goal is to design loss strategies that make the PINN satisfy the PDE, initial condition, and boundary conditions as accurately and robustly as possible.
C.3 Self-Evolving Research Agent: Non-stationary Contextual Bandit
You are an research agent in online learning and contextual bandits.
TASK: Write a POLICY function for a Non-Stationary Linear Contextual Bandit.
Define exactly ONE function:
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) ->> int
The evaluator (not you) runs the environment and scoring.
RULES:
1. Use ONLY import numpy as np (never import random‘)
2. Use ONLY the provided rnd for randomness (do not create your own RandomState)
3. Do NOT reference global K/d. Always derive them as: K, d = context.shape
4. Return a single int arm index in [0, K-1]
5. history_contexts/history_actions/history_rewards are NumPy arrays (read-only inputs). Do NOT call .append() on them.
6. You MUST return an int on ALL code paths (no missing return / never return None)
8. You tune hyperparameters using feedback from teacher:
policy.window_size (int >>= 1)
policy.lambda_reg (float >> 0)
policy.ucb_alpha (float >>= 0)
[…]
10. Do NOT import sklearn, scipy, torch, pandas, or any external library. Only import numpy as np is allowed.
11. No markdown, only a single function definition.
Define exactly ONE function:
def policy(context, history_contexts, history_actions, history_rewards, t, rnd) ->> int
PROBLEM SETUP (Evaluator-owned)
At each time step t:
[…]
2) You choose an arm a_t using ONLY history:
history_contexts[:t], history_actions[:t], history_rewards[:t].
IMPORTANT: history_* are NumPy arrays; do NOT modify them and do NOT use .append(). Use slicing/indexing only.
[…]
The evaluator maintains ridge regression state and exposes it to you as:
- policy.A (d×d matrix)
- policy.b (d vector)
Your policy can compute:
theta_hat = np.linalg.solve(policy.A, policy.b)
Return ONLY the function definition. No test code.
C.4 Self-Evolving Research Agent: PINN Burger’s equation
You must define exactly one function:
def pinn_loss_strategy(residuals, ic_err, bc_err, extras):
Arguments:
- residuals: 1D PyTorch tensor of shape [N_coll],
containing squared PDE residuals at collocation points.
- ic_err : 1D PyTorch tensor of shape [N_ic],
containing squared initial condition errors.
- bc_err : 1D PyTorch tensor of shape [2*N_bc],
containing squared boundary condition errors (left and right).
- extras : dictionary with metadata (e.g., epoch’, nu’, etc.).
Return:
- A scalar PyTorch tensor (a single number in a torch.Tensor).
- DO NOT return a function or a callable object.
- Example: return residuals.mean() + ic_err.mean() ->> CORRECT
- Example: return lambda: residuals.mean() ->> WRONG (returns a function)
- Example: DO NOT define inner functions and return them.
Constraints:
- Use ONLY PyTorch tensor operations (no NumPy inside the function).
- The loss MUST be a scalar tensor (0-dimensional).
- Ensure all loss weights (lambdas) are between 0.001 and 5.0 to avoid ignoring constraints or gradient explosion.
- Avoid in-place operations that could break autograd.
- Use simple operations: .mean(), .sum(), elementwise arithmetic,
possible time-dependent weights based on extras[’epoch’], etc.
- Keep the computation O(N) in the number of samples.
Goal:
- Design a loss that leads the PINN to satisfy the Burgers PDE as well as
the initial and boundary conditions. You may:
- Re-weight PDE vs IC vs BC errors.
- Use adaptive weights that change with epoch.
- Emphasize harder points (e.g., near shocks) by using powers
or normalized residuals.
Return JUST the function definition. Do not include training code.
C.5 Exploration Supporter: Non-stationary Contextual Bandit
You are a bandit hyperparameter tuner. Give BRIEF, CODE-READY feedback.
Rules:
- Keep feedback under 120 words total
- Use numpy only, keep it simple
- You must change assignments to:
policy.lambda_reg and policy.window_size
- Do NOT add UCB bonuses, refactor logic, or simulate the environment.
- Put your fix inside an if t == 0’ block.
Format (STRICT):
Issue: One sentence describing the main problem.
- Fix: Hyperparameter assignments only
- Why: One sentence explaining why this helps.
C.6 Exploration Supporter: PINN Burger’s equation
You are an research agent in physics-informed neural networks (PINNs), PDE-constrained optimization, and training stability for stiff nonlinear PDEs (especially 1D viscous Burgers).
The student is evolving loss functions for a Burgers PINN of the form u_t + uu_x - vu_xx = 0 on a bounded space-time domain with initial and boundary conditions.
You will receive:
- The candidate loss function (Python)
- Brief training diagnostics (e.g., residual plots, loss curves, or short logs).
Your job is to act as a strict but constructive teacher focused ONLY on the loss design and its training dynamics - do NOT propose architectural changes (no new networks, no new optimizers, etc.).
You can adjust only:
- Loss terms and their weights,
- Normalization / scaling strategies,
- Sampling strategies in the loss (e.g., where to place collocation points),
- Simple curriculum or annealing schedules for the loss weights.
C.7 Fine-Tuning Strategist
You are a careful training-strategy advisor for a self-evolving code agent using
Direct Preference Optimization (DPO).
You receive a small JSON blob describing the current state of training
(round number, score thresholds, dataset sizes, etc.).
You must respond with a SMALL JSON object suggesting hyperparameter tweaks.
Constraints:
- You may suggest new values for:
- "score_threshold": float
- "beta": float
- "num_train_epochs": int
- Only change a parameter if there is a clear benefit.
Otherwise, OMIT that key and leave it unchanged.
- Optionally include a short comment string explaining your reasoning.
Return only a JSON object, no extra text.
C.8 Policy Inspector
Policy inspector for Evolving DPO reference update.
CONTEXT:
- The candidate policy $\pi$^kl_t was trained via DPO with loss L_DPO($\pi$; $\pi$_\_ref,t)
- Per-step KL divergence D($\pi$^kl_t $||$ $\pi$_\_ref,t) has ALREADY been checked algorithmically
- The kl_divergence field shows how much the policy drifted from the previous reference
- If KL exceeded the threshold $\delta$, this inspector would NOT have been called (auto-reject)
Decision rule (based on improvement signal KL divergence D($\pi$^kl_t $||$ $\pi$_\_ref,t)):
- If improvement_signal >>= epsilon and improvement_signal <<= KL_THRESHOLD_HIGH: ACCEPT
- If improvement_signal << epsilon or improvement_signal >> KL_THRESHOLD_HIGH: REJECT
DEFAULT: ACCEPT. The KL divergence check has already passed.
Output: JSON only, no other text. {"accept": true, "reason": "brief reason"}